# Revision of two-temperature magnetically arrested flows onto a black hole

### v1

@@ Appendix A Light curves
The main text of this manuscript presents statistical properties of light-curves from MAD simulations with different spin values. Figures 13 through 15 present light-curves of individual zero and prograde models for the default viewing angle $i=160$ deg. Here in addition to 86 and 229 GHz we also show light-curves at 690 GHz for visual comparison.
Figure 13: Radiative signatures for MAD $a_{*}=0$ model at the default viewing angle of $i=160$ deg. Panels from left to right display Stokes ${\mathcal{I}}$ (86, 229, 690 GHz), RM (213-229 GHz), LP (86, 229, 690 GHz) and CP (86, 229, 690 GHz). Panels from top to bottom show models R1-160, and K.
@@ Appendix B GRRMHD
In ebhlight the multifrequency radiative transport is carried out using Monte Carlo scheme in which the radiation field is simulated with a large number of superphoton particles representing even larger number of physical photons (see code description in Ryan et al. (2015) for exact definitions). To make sure that the radiative cooling of plasma is well captured, for each fluid element the Monte Carlo scheme has to produce a sufficient number of photon particles within the electron cooling timescale.
The electron cooling timescale (in M units) writes:
@@
where $u_{\rm e,code}$ is the electron internal energy provided by the two-temperature model (in the code units) and where $\Lambda_{\rm code}$ is the synchrotron cooling rate (also in the code units). In c.g.s units the cooling rate reads (Eq. A4 from Mościbrodzka et al. 2011):
@@
where conversion to code units is: $\Lambda_{\rm code}=\Lambda_{\rm cgs}{\mathcal{L}}{\mathcal{T}^{3}}/{\mathcal{M}}$. Figure 16 shows GRRMHD quality factor $Q_{\rm em}$, a number of superphotons emitted within $\tau_{\rm cool}$, for a snapshot of GRRMHD around $t$=10, [^0] M. $Q_{\rm em}\gg 10$ in the inner disk ($r<10M$) and along the jet wall, indicating reasonable sampling in regions where the radiative cooling is the most efficient.
@@
Appendix C The issue of $R_{\rm low}$ parameter in $R(\beta)$ model and parameter inference
When estimating electron temperature parameters using polarimetric images of Sgr A*, Event Horizon Telescope Collaboration et al. (2024) assumed that in Equation 3, $R_{\rm low}=1$. What happens if we relax the assumption about $R_{\rm low}$? Figure 17 shows the impact of varying $R_{\rm low}$ parameter on the linear and circular polarimetric images of Sgr A*. Images of the MAD with ($R_{\rm low},R_{\rm high})=(10,10)$ start to resemble those with ($R_{\rm low},R_{\rm high})=(1,160)$. It is therefore possible that polarimetric scoring of models with $R_{\rm low}\sim 10$ could lead to different best-bet values of $R_{\rm high}\ll 160$. If this is indeed the case, which has to be carefully verified in the future, electron temperature parameter estimations using $R(\beta)$ models with $R_{\rm low}\sim 10$ have a better chance to converge with some of the discussed $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ models, proving that particle-in-cell models applied to GRMHD simulations correctly predict the order of magnitude of electron heating in the collisionless plasma.
Figure 17: Polarimetric images of a single GRMHD MAD $a_{*}=0.5$ snapshot. Panels show the same quantities as in Figure 6 but only for $R(\beta)$ models with different assumptions of $R_{\rm low}$ parameter in Equation 3. All models are normalized with different ${\mathcal{M}}$ to reproduce the same total flux. From left to right: ${\mathcal{M}}/10^{17}$=3,4,5,8.

### v3

@@ Appendix A Light curves
The main text of this manuscript presents statistical properties of light curves from MAD simulations with different spin values. Figures 13 - 15 present light curves of individual zero and prograde models for the default viewing angle $i=160^{\circ}$. Here in addition to 86 and 229 GHz we also show light curves at 690 GHz for visual comparison.
Figure 13: Radiative signatures for the MAD $a_{*}=0$ model at the default viewing angle of $i=160^{\circ}$. Panels from left to right display Stokes ${\mathcal{I}}$ (86, 229, 690 GHz), RM (213 - 229 GHz), LP (86, 229, 690 GHz) and CP (86, 229, 690 GHz). Panels from top to bottom show models R1-160, and K.
@@ Appendix B GRRMHD
In ebhlight, the multifrequency radiative transport is carried out using the Monte Carlo scheme in which the radiation field is simulated with a large number of superphoton particles representing an even larger number of physical photons (see code description in Ryan et al. 2015 for exact definitions). To make sure that the radiative cooling of plasma is well captured, for each fluid element, the Monte Carlo scheme has to produce a sufficient number of photon particles within the electron cooling timescale.
The electron cooling timescale (in M units) can be written as
@@
where $u_{\rm e,code}$ is the electron internal energy provided by the two-temperature model (in code units) and where $\Lambda_{\rm code}$ is the synchrotron cooling rate (also in code units). In cgs units the cooling rate reads (Eq. (A4) from Mościbrodzka et al. 2011)
@@
where the conversion to code units is $\Lambda_{\rm code}=\Lambda_{\rm cgs}{\mathcal{L}}{\mathcal{T}^{3}}/{\mathcal{M}}$. Figure 16 shows the quality factor of GRRMHD $Q_{\rm em}$, a number of superphotons emitted within $\tau_{\rm cool}$, for a snapshot of GRRMHD around $t$ = 10, [^0] M. $Q_{\rm em}\gg 10$ in the inner disk ($r<10$M) and along the jet wall, indicating reasonable sampling in regions where radiative cooling is the most efficient.
@@
Appendix C The issue of $R_{\rm low}$ parameter in $R~{}(\beta)$ model and parameter inference
When estimating the parameters of the electron temperature using polarimetric images of Sgr A*, Event Horizon Telescope Collaboration et al. (2024) assumed that in Equation 3, $R_{\rm low}=1$. What happens if we relax the assumption about $R_{\rm low}$? Figure 17 shows the impact of the varying parameter $R_{\rm low}$ on the linear and circular polarimetric images of Sgr A*. The images of the MAD with ($R_{\rm low},R_{\rm high})=(10,10)$ start to resemble those of ($R_{\rm low},R_{\rm high})=(1,160)$. Therefore, it is possible that the polarimetric scoring of the models with $R_{\rm low}\sim 10$ could lead to different best-bet values of $R_{\rm high}\ll 160$. If this is indeed the case, which will have to be carefully verified in the future, estimations of the electron temperature parameters using $R~{}(\beta)$ models with $R_{\rm low}\sim 10$ have a better chance of convergence with some of the discussed $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ models, proving that particle-in-cell models applied to GRMHD simulations correctly predict the order of magnitude of electron heating in collisionless plasma.
Figure 17: Polarimetric images of a single GRMHD MAD $a_{*}=0.5$ snapshot. Panels show the same quantities as in Figure 6 but only for $R~{}(\beta)$ models with different assumptions of $R_{\rm low}$ parameter in Equation 3. All models are normalized with different ${\mathcal{M}}$ to reproduce the same total flux. From left to right: ${\mathcal{M}}/10^{17}$ = 3, 4, 5, 8.